OpenStudy (anonymous):

Is the equation 11-3x^2+7x-4=8x+1 equivalent to 3x^2+x-6=0 ???

OpenStudy (bahrom7893):

simplify the first equation. What do you get?

OpenStudy (anonymous):

3x^2+x-6=0 Im just double checking because I always mess up somewhere when I simplify the equations so I just like to double check

OpenStudy (bahrom7893):

11-3x^2+7x-4=8x+1 -3x^2 + 7x - 8x - 1 + 11 - 4 = 0 -3x^2 - x + 6 = 0

OpenStudy (bahrom7893):

Multiply everything out by -1 to get: 3x^2 + x - 6 = 0

OpenStudy (bahrom7893):

Yes they are

OpenStudy (anonymous):

Woo hoo! I finally did it right!! thank you!